# Unified approach to the resources of tensor network and stabilizer simulations

### v2

@@ Abstract
Operator stabilizer formalism.—When $|\psi\rangle=U_{c}|0\rangle^{\otimes n}$ with $U_{c}$ a Clifford circuit, its simulation can be efficiently done classically through the stabilizer formalism. The key idea is that the initial state $|0\rangle^{\otimes n}$ has $n$ stabilizers $\{Z_{i}\}$: $|0\rangle\langle 0|^{\otimes n}=\prod_{i}(I_{n}+Z_{i})/2$ and the state $|\psi\rangle$ has also $n$ $n$-qubit stabilizers $\{U_{c}Z_{i}U_{c}^{\dagger}\}$: $|\psi\rangle\langle\psi|=\prod_{i}\left(I_{n}+U_{c}Z_{i}U_{c}^{\dagger}\right)/2$. Since $U_{c}Z_{i}U_{c}^{\dagger}$ are still Pauli operators, we are able to efficiently track them and further simulate $|\psi\rangle$ classically in both weak and strong ways [2]. For general $U$ with magic gates ($T$) in it, $U_{c}Z_{i}U_{c}^{\dagger}$ will become a linear combination of many Pauli operators, which increases the complexity of classical simulation with exponential dependence on the number $N_{T}$ of $T$ gates [2, 24].
The operator stabilizer formalism (OSF) is a generalization of standard stabilizer formalism for simulating $UOU^{\dagger}$ with arbitrary $O$. The idea is straightforward in the vectorization space. When $|O\rangle$ itself is a stabilizer state in vectorization space with $2n$ $2n$-qubit Pauli stabilizers $\{S_{O,i}\}$ and $U=U_{c}$, we have
$U_{c}\otimes U_{c}^{*}|O\rangle\langle O|U_{c}^{\dagger}\otimes U_{c}^{T}$
$=\prod_{i}\frac{I_{2n}+(U_{c}\otimes U_{c}^{*})S_{O,i}(U_{c}^{\dagger}\otimes U_{c}^{T})}{2}.$ (17)
Since $U_{c}\otimes U_{c}^{*}$ is still Clifford, stabilizers of $U\otimes U^{*}|O\rangle$ are still Pauli operators, we are able to efficiently track them and further simulate $|\psi\rangle$ classically such as estimating operator overlap $\mathrm{Tr}(O^{\prime{\dagger}}U_{c}OU_{c}^{\dagger})=\langle O^{\prime}|U_{c}\otimes U_{c}^{*}|O\rangle$ when $|O^{\prime}\rangle$ is also a stabilizer state [2].
The classical complexity of OSF in vectorization space also depends on magic. Here, magic can appear in two places. First is the magic contained in $|O\rangle$ since not every operator $O$ has a stabilizer vectorization state. Examples with zero magic are when $O=|i\rangle\langle j|$ is a matrix element basis and when $O=2^{-n/2}P_{i}$ is a Pauli operator. For other $|O\rangle$, we can quantify its magic by the minimal number $N_{O,T}$ of $T$ gates that is needed to prepared $|O\rangle$ from a stabilizer state. Second is the magic contained in $U\otimes U^{*}$ which can be further characterized by $N_{T}$. We can summarize this in the following theorem.
Theorem 4.
The complexity of classically simulating $UOU^{\dagger}$ through OSF is $\mathcal{O}\left(\exp(N_{T}+N_{O,T})\right)$.
While just like the standard stabilizer formalism, the complexity of OSF generally exhibits an exponential dependence on $N_{T}$, there are cases with many $T$ gates where directly using the standard stabilizer formalism can be very expensive, while OSF is efficient. The idea is that, although $T$ gates can create Pauli superposition (or BBC), $|O\rangle$ with large BBC may still be a stabilizer state in vectorization space. This can be seen from the action of $T$: $TXT^{\dagger}=\frac{1}{\sqrt{2}}(X+Y),TYT^{\dagger}=\frac{1}{\sqrt{2}}(-X+Y)$ [45, 46], which, while creating Pauli superpositions, effectively acts as a Hadamard gate in the subspace $\{|X\rangle,|Y\rangle\}$. We give two such examples in SM F. One is the Magic-IQP circuit with $U=T^{\otimes n}U_{c}T^{\otimes n}$ and the other is the $X$-$Y$ preserving circuit with $U$ is composed of $\{T,SWAP,C_{3}\}$ with $C_{3}$ a special three-qubit Clifford gate, as shown in Ref [45].
Figure 3: Relations between tensor network and stabilizer formalism. As BKE increases, SE, which governs the complexity of the tensor network approach, will change from mainly decided by CBC ($H$) to mainly decided by BBC ($T$). On the other hand, OSF is always governed by magic $T$. Therefore, as BKE increases, the related resources between the two approaches will go from mismatching to matching.
Resource connections between tensor network and stabilizer formalism.— In the tensor network approach, SE (bond dimension) determines the computational complexity. When the BKE of $|O\rangle$ is low, SE primarily depends on the CBC resource in $U$, which is related to creating state computational basis superposition and is the key for the separation between quantum computing and reversible classical computing [47], often measured by the number of Hadamard gates. In contrast, when BKE is high, SE is mainly governed by the magic resource linked to BBC, associated with Pauli superpositions and typically measured by the number of $T$ gates. In the stabilizer formalism, however, only the magic resource (i.e., $T$ count) contributes to complexity growth. Therefore, as BKE increases, the dominant resources in both approaches shift from unrelated to strongly correlated. This key transition is summarized in Fig. 3. Nevertheless, exceptions exist, for example, X–Y preserving circuits [45] can yield large SE while remaining efficiently simulatable by OSF.
Discussion and outlooks.—Tensor network methods and the stabilizer formalism are two principal approaches for classically simulating quantum circuits. While both have been extensively studied, their relation remains poorly understood. In this work, we give a unified understanding through vectorization and show that the bra-ket entanglement of the operator highly influences the correlation level between the two approaches.
@@
Several open directions remain. Extending our framework to open quantum systems seems natural but challenging: quantum channels are merely non-unitary gates in the vectorization space, but can alter BKE. Also, although the computational basis $|\ket{i}\bra{j}$ and the Pauli basis $P_{i}$ can be viewed as gauge choices and can be unitarily transformed to other bases with the same BKE, they correspond to extreme minimal and maximal BKE values, respectively. This raises the question of whether such extreme bases possess special positions, and whether results such as Theorems 12 can be generalized to arbitrary matrix bases.
@@ B.1 Space entanglement (SE)
$H_{SE,\alpha}\left(|O\rangle\right)=\frac{1}{1-\alpha}\log\left(\sum_{i}s_{i}^{2\alpha}\right).$ (28)
@@
$H_{BKE,\alpha}(|O\rangle)=\frac{1}{1-\alpha}\log\left(\sum_{i}\gamma_{i}^{2\alpha}\right).$ (30)
@@
$H_{FBKE,\alpha}(|O\rangle)=\frac{1}{1-\alpha}\log\left(\sum_{i}\lambda_{i}^{2\alpha}\right).$ (34)
@@
$H_{BKE,\alpha}(|O\rangle)+H_{FBKE,\beta}(|O\rangle)\geq n,\quad\forall~\alpha,\beta\geq\frac{1}{2}\quad\text{and}\quad\frac{1}{\alpha}+\frac{1}{\beta}=2.$ (35)
@@
$H_{BKE,\alpha}\left(|O\rangle\right)$ $=H_{BKE,\alpha}\left(\mathcal{V}\left[UOU^{\dagger}\right]\right),$ (37)
$H_{FBKE,\alpha}\left(|O\rangle\right)$ $=H_{FBKE,\alpha}\left(\mathcal{V}\left[UOU^{\dagger}\right]\right)$ (38)
@@ C.1 Computational basis coherence (CBC)
$H_{CBC,\alpha}(|O\rangle)=\frac{1}{1-\alpha}\log\left(\sum_{ij}|o_{ij}|^{2\alpha}\right).$ (40)
For example, when $O=|i\rangle\langle j|$ is a matrix element basis, we have $H_{CBC,\alpha}(|i\rangle|j\rangle)=0$. In contrast, when $O=2^{-n/2}P_{i}$ is a Pauli basis, we have $H_{CBC,\alpha}(|B_{i}\rangle)=n$.
@@
$H_{CBC,\alpha}\left(U\otimes U^{*}|O\rangle\right)\leq\min\{2N_{H}+H_{CBC,0}(|O\rangle),2n\}.$ (41)
@@
$H_{CBC,\alpha}\left(U\otimes U^{*}|O\rangle\right)\leq H_{CBC,0}\left(U\otimes U^{*}|O\rangle\right).$ (42)
@@
$H_{CBC,0}\left(U\otimes U^{*}|O\rangle\right)$ $\leq N_{H}\times\log 4+H_{CBC,0}(|O\rangle)$ (44)
$=2N_{H}+H_{CBC,0}(|O\rangle).$
Also, since there are $2^{2n}$ computational basis in vectorization space, another upper bound of $H_{CBC,0}\left(U\otimes U^{*}|O\rangle\right)$ is $2n$. ∎
@@ C.2 Bell basis coherence (BBC)
$H_{BBC,\alpha}(|O\rangle)=\frac{1}{1-\alpha}\log\left(\sum_{i}|b_{i}|^{2\alpha}\right).$ (46)
For example, when $O=|i\rangle\langle j|$ is a matrix element basis, we have $H_{BBC,\alpha}(|i\rangle|j\rangle)=n$. In contrast, when $O=2^{-n/2}P_{i}$ is a Pauli basis, we have $H_{BBC,\alpha}(|B_{i}\rangle)=0$.
BBC has strong connections with magic. We can consider the changing of BBC between $|O\rangle$ and $U\otimes U^{*}|O\rangle$. Since BBC characterizes the distribution of $O$ on the Pauli basis, when $U$ is constructed from the standard gate set $\{H,S,T,CX\}$, the $T$ gate is the only gate that can change $H_{BBC,\alpha}(|O\rangle)$ since the other gates are all Clifford which are permutation gates on Pauli basis. The effect of $T$ on BBC can be summarized in the following lemma.
@@
$H_{BBC,\alpha}\left(U\otimes U^{*}|O\rangle\right)\leq\min\{N_{T}+H_{BBC,0}(|O\rangle),2n\}.$ (47)
@@
$H_{BBC,\alpha}\left(U\otimes U^{*}|O\rangle\right)\leq H_{BBC,0}\left(U\otimes U^{*}|O\rangle\right).$ (48)
@@
$H_{BBC,0}\left(U\otimes U^{*}|O\rangle\right)\leq N_{T}+H_{BBC,0}(|O\rangle).$ (49)
Also, since there are $2^{2n}$ Bell basis in vectorization space, another upper bound of $H_{BBC,0}\left(U\otimes U^{*}|O\rangle\right)$ is $2n$. ∎
@@ D.1 CBC and BBC bound SE
[…]
@@
$H_{\alpha}\left(\{|c_{\alpha\beta}|^{2}\}\right)\geq H_{SE,\alpha}(|O\rangle).$ (54)
@@
[Theorem 1 in main text] $G_{zero,\alpha}(|O\rangle)$ has the upper bound
$G_{zero,\alpha}(|O\rangle)\leq\min\left\{n,H_{CBC,\alpha}(|O\rangle),H_{BBC,\alpha}(|O\rangle)\right\}-H_{SE,\alpha}(|O\rangle).$ (55)
$G_{CBC,\alpha}(|O\rangle)$ has the upper bound
$G_{CBC,\alpha}(|O\rangle)\leq\min\left\{n,H_{CBC,\alpha}(|O\rangle)\right\}-H_{SE,\alpha}(|O\rangle).$ (56)
$G_{BBC,\alpha}(|O\rangle)$ has the upper bound
$G_{BBC,\alpha}(|O\rangle)\leq\min\left\{H_{BBC,\alpha}(|O\rangle),n\right\}-H_{SE,\alpha}(|O\rangle).$ (57)
@@
$H_{SE,\alpha}\left(U\otimes U^{*}|O\rangle\right)$ $\leq n$ (58)
$H_{SE,\alpha}\left(U\otimes U^{*}|O\rangle\right)$ $\leq H_{CBC,\alpha}(|O\rangle)$ (59)
$H_{SE,\alpha}\left(U\otimes U^{*}|O\rangle\right)$ $\leq H_{BBC,\alpha}(|O\rangle)$ (60)
Since we have
$H_{SE,\alpha}\left(U\otimes U^{*}|O\rangle\right)\leq H_{SE,0}\left(U\otimes U^{*}|O\rangle\right)\leq n,$ (61)
$H_{CBC,\alpha}(|O\rangle)=H_{CBC,\alpha}\left(U\otimes U^{*}|O\rangle\right),$ (62)
$H_{BBC,\alpha}(|O\rangle)=H_{BBC,\alpha}\left(U\otimes U^{*}|O\rangle\right).$ (63)
[…]
$H_{CBC,\alpha}(|O\rangle)\geq H_{SE,\alpha}(|O\rangle),$ (64)
$H_{BBC,\alpha}(|O\rangle)\geq H_{SE,\alpha}(|O\rangle).$ (65)
@@
$H_{CBC,\alpha}(|O\rangle)$ $=H_{\alpha}\left(\{|c_{\alpha\beta}|^{2}\}\right)\geq H_{SE,\alpha}(|O\rangle),$ (66)
$H_{BBC,\alpha}(|O\rangle)$ $=H_{\alpha}\left(\{|b_{j}|^{2}\}\right)\geq H_{SE,\alpha}(|O\rangle).$ (67)
Therefore, when no additional CBC and BBC resources ($H$ and $T$) are introduced, $G_{zero,\alpha}(|O\rangle)$ is governed by the existing CBC and BBC in $|O\rangle$. To generate higher entanglement, additional CBC and BBC resources are needed. Since CBC and BBC are about projecting $|O\rangle$ on two different bases, following Lemma 1, we can summarize a CBC-BBC uncertainty principle.
@@
$H_{CBC,\alpha}(|O\rangle)+H_{BBC,\beta}(|O\rangle)\geq n,\quad\forall~\alpha,\beta\geq\frac{1}{2}\quad\text{and}\quad\frac{1}{\alpha}+\frac{1}{\beta}=2.$ (68)

### v3

@@ Abstract
Implications for classical simulation.—We now show three implications of our results on classical simulations. The first two focus on the resource dependence transition for simulation methods related to entanglement, such as matrix product states and operators [35, 16]. The last one focus on the relations between different types of simulation methods.
Resource transition in marginal probability simulation. BKE-matching operators arise naturally in practical quantum simulation tasks (see Appendix F for details and other generalizations). As a key example, computing marginal probability distributions [9] over $n_{2}$ qubits of an $n=(n_{1}+n_{2})$-qubit circuit $U$ amounts to evaluating
$p(\mathbf{s})=\mathrm{tr}[|0\rangle\!\langle 0|^{\otimes n}U^{\dagger}(\mathbb{I}_{n_{1}}\otimes|\mathbf{s}\rangle\!\langle\mathbf{s}|U)].$ (18)
The effective observable $O=\mathbb{I}_{n_{1}}\otimes|\mathbf{s}\rangle\!\langle\mathbf{s}|$ is a BKE-matching operator with $r=n_{1}$, so Corollary 1 directly yields $N_{T}\geq\eta-n_{2}$ and $N_{H}\geq\eta-n_{1}$. This interpolates between the coherence-limited regime for amplitude estimation ($n_{1}=0$) and the magic-limited regime for Pauli average estimation ($n_{1}=n-1$)—two tasks that, even for the same $T$-depth-one circuit, exhibit a sharp complexity separation from GapP-complete to P [49].
Resource dependence transition in mixed state simulation. When $O=\rho$ is a density matrix, the order-2 BKE is then related to the purity of the state
$H_{\mathrm{BKE},2}(\rho)=-\log\mathrm{tr}(\rho^{2}).$ (19)
Thus, using our estabilished picture above, we have the story for mixed state simulation. As purity decreases, BKE increases, shifting the dominant resource for SE growth from coherence ($N_{H}$) to magic ($N_{T}$). This implies that high-temperature quantum simulations are more sensitive to magic, while low-temperature ones are more sensitive to coherence.
Relating different classical simulation methods. There are classical simulation methods that are directly related to entanglement, coherence, and magic, respectively. Examples are tensor-network-based methods (entanglement) [34, 16, 35], Stabilizer and Pauli truncation-based methods (magic) [24, 1, 39, 2, 4, 5], and Feynman path-based methods (coherence) [12, 18]. Since Theorem 1, can be understood as BBC and CBC are lower bounded by SE, therefore, whenever a quantum system $UOU^{\dagger}$ is difficult for entanglement-related methods, it must also be difficult for magic and coherence-related methods. Also, when BKE is low, entanglement-related methods can be more similar to coherence-related methods in terms of the algorithm complexity. In contrast, when BKE is high, entanglement-related methods will be more approaching to magic-related methods.
Outlooks.—The desire to connect different concepts is rooted in the genes of physics. In this work, by introducing bra-ket entanglement as a diagnostic quantity, we have successfully bridged entanglement, coherence, and magic in a unified picture. We anticipate our results to have both significant impacts in quantum resource theories and in understanding the boundary between classical simulatabilities and quantum advantages.
@@
Several open directions remain. Extending our framework to open quantum systems seems natural but challenging: quantum channels are merely non-unitary gates in the vectorization space, but can alter BKE. Also, although the computational basis and the Pauli basis can be viewed as gauge choices and can be unitarily transformed to other bases with the same BKE, they correspond to extreme minimal and maximal BKE values, respectively. This raises the question of whether such extreme bases possess special positions, and whether results such as Theorems 2 can be generalized to arbitrary matrix bases. Finally, since both CBC and BBC have non-zero lower bounds and therefore non-zero upper bounds for SE, no existing simulation methods have efficient and bounded complexity even for the circuit family $\mathcal{F}_{zero}$. Thus, we suggest for exploring specially-designed simulation methods for operators with intermediate BKE.
Code: We apply Mindspore Quantum [46] for our numerical simulations with the source code [13].
@@ B.1 Space entanglement (SE)
$H_{SE,\alpha}(O)=\frac{1}{1-\alpha}\log\left(\sum_{i}s_{i}^{2\alpha}\right).$ (30)
@@
$H_{BKE,\alpha}(O)=\frac{1}{1-\alpha}\log\left(\sum_{i}\gamma_{i}^{2\alpha}\right).$ (32)
@@
$H_{FBKE,\alpha}(O)=\frac{1}{1-\alpha}\log\left(\sum_{i}\lambda_{i}^{2\alpha}\right).$ (36)
@@
$H_{BKE,\alpha}(O)+H_{FBKE,\beta}(O)\geq n,\quad\forall~\alpha,\beta\geq\frac{1}{2}\quad\text{and}\quad\frac{1}{\alpha}+\frac{1}{\beta}=2.$ (37)
@@
$H_{BKE,\alpha}(O)$ $=H_{BKE,\alpha}\left(UOU^{\dagger}\right),$ (39)
$H_{FBKE,\alpha}(O)$ $=H_{FBKE,\alpha}\left(UOU^{\dagger}\right)$ (40)
@@ C.1 Computational basis coherence (CBC)
$H_{\mathrm{CBC},\alpha}(O)=\frac{1}{1-\alpha}\log\left(\sum_{ij}|o_{ij}|^{2\alpha}\right).$ (42)
For example, when $O=|i\rangle\langle j|$ is a matrix element basis, we have $H_{\mathrm{CBC},\alpha}(|i\rangle\!\langle j|)=0$. In contrast, when $O=2^{-n/2}P_{i}$ is a Pauli basis, we have $H_{\mathrm{CBC},\alpha}(2^{-n/2}P_{i})=(1-\alpha)^{-1}\log\left[2^{-\alpha n}\cdot 2^{n}\right]=n$.
@@
$H_{\mathrm{CBC},\alpha}(U^{\dagger}OU)\leq\min\{2N_{H}+H_{CBC,0}(O),2n\}.$ (43)
@@
$H_{\mathrm{CBC},\alpha}(U^{\dagger}OU)\leq H_{CBC,0}(U^{\dagger}OU).$ (44)
@@
$H_{CBC,0}(U^{\dagger}OU)$ $\leq N_{H}\times\log 4+H_{CBC,0}(O)$ (46)
$=2N_{H}+H_{CBC,0}(O).$
Also, since there are $2^{2n}$ computational basis in vectorization space, another upper bound of $H_{CBC,0}(U^{\dagger}OU)$ is $2n$. ∎
@@ C.2 Bell basis coherence (BBC)
$H_{\mathrm{BBC},\alpha}(O)=\frac{1}{1-\alpha}\log\left(\sum_{i}|b_{i}|^{2\alpha}\right).$ (48)
For example, when $O=|i\rangle\langle j|$ is a matrix element basis, we have $H_{\mathrm{BBC},\alpha}(|i\rangle|j\rangle)=n$. In contrast, when $O=2^{-n/2}P_{i}$ is a Pauli basis, we have $H_{\mathrm{BBC},\alpha}(|B_{i}\rangle)=0$.
BBC has strong connections with magic. We can consider the changing of BBC between $|O\rangle$ and $U\otimes U^{*}|O\rangle$. Since BBC characterizes the distribution of $O$ on the Pauli basis, when $U$ is constructed from the standard gate set $\{H,S,T,CX\}$, the $T$ gate is the only gate that can change $H_{\mathrm{BBC},\alpha}(O)$ since the other gates are all Clifford which are permutation gates on Pauli basis. The effect of $T$ on BBC can be summarized in the following lemma.
@@
$H_{\mathrm{BBC},\alpha}(U^{\dagger}OU)\leq\min\{N_{T}+H_{BBC,0}(O),2n\}.$ (49)
@@
$H_{\mathrm{BBC},\alpha}(U^{\dagger}OU)\leq H_{BBC,0}(U^{\dagger}OU).$ (50)
@@
$H_{BBC,0}(U^{\dagger}OU)\leq N_{T}+H_{BBC,0}(O).$ (51)
Also, since there are $2^{2n}$ Bell basis in vectorization space, another upper bound of $H_{BBC,0}(U^{\dagger}OU)$ is $2n$. ∎
@@ D.1 CBC and BBC bound SE
Given a $n$ qubit system with basis $\{B_{1},\cdots,B_{n^{2}}\}$ such that $\text{tr}(B_{j}B_{k}^{\dagger})=\delta_{jk}$ and $B_{j}=O_{j}^{(1)}\otimes\cdots\otimes O_{j}^{(n)}$ are tensor product operators. Then the decomposition $O=\sum_{j=1}^{n^{2}}c_{j}B_{j}$ satisfied $H_{\alpha}(\{c_{i}^{2}\})\geq H_{SE,\alpha}(O)$
@@
$H_{\alpha}\left(\{|c_{\alpha\beta}|^{2}\}\right)\geq H_{SE,\alpha}(O).$ (56)
@@
[Theorem 1 in main text] $G_{zero,\alpha}(O)$ has the upper bound
$G_{zero,\alpha}(O)\leq\min\left\{n,H_{\mathrm{CBC},\alpha}(O),H_{\mathrm{BBC},\alpha}(O)\right\}-H_{SE,\alpha}(O).$ (57)
$G_{\mathrm{CBC},\alpha}(O)$ has the upper bound
$G_{\mathrm{CBC},\alpha}(O)\leq\min\left\{n,H_{\mathrm{CBC},\alpha}(O)\right\}-H_{SE,\alpha}(O).$ (58)
$G_{\mathrm{BBC},\alpha}(O)$ has the upper bound
$G_{\mathrm{BBC},\alpha}(O)\leq\min\left\{H_{\mathrm{BBC},\alpha}(O),n\right\}-H_{SE,\alpha}(O).$ (59)
@@
Because Eq. 58-59 are just corollaries of Eq. 57, it is sufficient to prove the first equation. By the definition of $G_{\mathcal{F},\alpha}(O)$, it is sufficient to show that $\forall U\in\mathcal{F}_{\mathrm{zero}}$,
$H_{SE,\alpha}(UOU^{\dagger})$ $\leq n$ (60)
$H_{SE,\alpha}(UOU^{\dagger})$ $\leq H_{\mathrm{CBC},\alpha}(O)$ (61)
$H_{SE,\alpha}(UOU^{\dagger})$ $\leq H_{\mathrm{BBC},\alpha}(O)$ (62)
Since we have
$H_{SE,\alpha}(UOU^{\dagger})\leq H_{SE,0}(UOU^{\dagger})\leq n,$ (63)
$H_{\mathrm{CBC},\alpha}(O)=H_{\mathrm{CBC},\alpha}(UOU^{\dagger}),$ (64)
$H_{\mathrm{BBC},\alpha}(O)=H_{\mathrm{BBC},\alpha}(UOU^{\dagger}).$ (65)
@@
$H_{\mathrm{CBC},\alpha}(O)\geq H_{SE,\alpha}(O),$ (66)
$H_{\mathrm{BBC},\alpha}(O)\geq H_{SE,\alpha}(O).$ (67)
@@
$H_{\mathrm{CBC},\alpha}(O)$ $=H_{\alpha}\left(\{|c_{\alpha\beta}|^{2}\}\right)\geq H_{SE,\alpha}(O),$ (68)
$H_{\mathrm{BBC},\alpha}(O)$ $=H_{\alpha}\left(\{|b_{j}|^{2}\}\right)\geq H_{SE,\alpha}(O).$ (69)
Therefore, when no additional CBC and BBC resources ($H$ and $T$) are introduced, $G_{zero,\alpha}(O)$ is governed by the existing CBC and BBC in $|O\rangle$. To generate higher entanglement, additional CBC and BBC resources are needed. Since CBC and BBC are about projecting $|O\rangle$ on two different bases, following Lemma 2, we can summarize a CBC-BBC uncertainty principle.
@@
$H_{\mathrm{CBC},\alpha}(O)+H_{BBC,\beta}(O)\geq n,\quad\forall~\alpha,\beta\geq\frac{1}{2}\quad\text{and}\quad\frac{1}{\alpha}+\frac{1}{\beta}=2.$ (70)